# The $S_{n}$-equivariant Euler characteristic of ¯$\overline{\mathcal{M}}_{1,n}(\mathbb{P}^{r},d)$

### v1

@@ 1.3. Enumerative combinatorics of cycles
$\phantom{space}+\sum_{d\geq 2}q^{d}\sum_{k=2}^{d}\eta_{k,d}(r)\frac{(1+2\dot{\mathscr{A}}_{0})^{2}}{(1-\psi_{2}(\mathscr{A}_{0}^{\prime\prime}))^{k/2+1}}+\sum_{j\mid k}\theta_{j,k,d}(r)\frac{1}{(1-\psi_{j}(\mathscr{A}_{0}^{\prime\prime}))^{k/j}}.$
[…]
@@
1.4. A further application
This paper is the first step of a larger programme to study motives of torus-fixed stable maps. In §7, we explain how our perspective yields the following structural result on the $S_{n}$-equivariant Serre characteristic in arbitrary genus.
Theorem C.
There exists a polynomial $\mathsf{K}_{g,n,d}(t)\in(K_{0}(\mathsf{M})\otimes\Lambda)[t]$ of degree $d$ in $t$ such that
$\mathsf{e}^{S_{n}}(\overline{\mathcal{M}}_{g,n}(\mathbb{P}^{r},d)^{\mathbb{C}^{\star}})=(r+1)\cdot\mathsf{K}_{g,n,d}(r+1)$
for all integers $r>0$.
We view Theorem C as a consequence of basic properties of chromatic polynomials of unlabeled graphs, as studied by Hanlon [Han85]. It has the following application to the topological Euler characteristic. For a partition $\lambda\vdash n$, we set
(1.7) $\chi_{\lambda}(\overline{\mathcal{M}}_{g,n}(\mathbb{P}^{r},d)):=\langle s_{\lambda},\chi^{S_{n}}(\overline{\mathcal{M}}_{g,n}(\mathbb{P}^{r},d))\rangle\in\mathbb{Z}.$
Above, $s_{\lambda}\in\Lambda$ is the Schur function corresponding to $\lambda$, and the brackets denote the Hall inner product on $\Lambda$. The statistic $\chi_{\lambda}(\overline{\mathcal{M}}_{g,n}(\mathbb{P}^{r},d))$ is the multiplicity, in the Euler characteristic of ¯$\overline{\mathcal{M}}_{g,n}(\mathbb{P}^{r},d)$, of the irreducible $S_{n}$-representation corresponding to $\lambda$. With this notation, Theorem C implies that there exist unique integers $a_{g,d}^{\lambda}(1),\ldots,a_{g,d}^{\lambda}(d+1)\in\mathbb{Z}$ such that for all target dimensions $r>0$,
$\chi_{\lambda}(\overline{\mathcal{M}}_{g,n}(\mathbb{P}^{r},d))=\sum_{k=1}^{d+1}a_{g,d}^{\lambda}(k)\binom{r+1}{k}.$
See Corollary 7.2. This pattern can be observed for $g=1$ in Tables 1, 2, and 3.
A basic corollary of Theorem C is that the $S_{n}$-equivariant Serre characteristic of ¯$\overline{\mathcal{M}}_{g,n}(\mathbb{P}^{r},d)^{\mathbb{C}^{\star}}$ is determined, for arbitrary $r$, by the values it takes on the first $d$ values of $r$. This seems to have interesting connections to the geometry of linear systems on nodal curves, and points to stablisation phenomena that we hope to address in future work. See §7 for further discussion of Theorem C and its consequences, as well as low degree calculations produced from our main formulas.
@@ 1.5. History and context
We are inspired by work of Getzler [Get95a, Get98] and Getzler–Kapranov [GK98] from the 1990s, using symmetric functions to work systematically with $S_{n}$-equivariant graph stratifications of moduli spaces of stable curves. Getzler and Kapranov work in the setting of modular operads, and give a formula [GK98, Theorem 8.13] which encodes how the two generating functions
@@
in $K_{0}(\mathsf{M})\otimes\Lambda[[q]]$ determine one another, in terms of plethysm and other algebraic operations on $\Lambda[[q]]$. Getzler computed $\mathsf{m}$ for $g\leq 1$ geometrically, and he then used the symmetric function formalism to compute $\overline{\mathsf{m}}$ for $g\leq 1$ [Get95a, Get98, Get99]. In higher genus, the framework developed by Getzler and Kapranov has become a standard tool for passing cohomological information between moduli spaces of smooth and stable curves: see [BT07, §2] or [PW21, §2.5]. Gorsky has computed the $S_{n}$-equivariant Euler characteristic $\chi^{S_{n}}(\mathcal{M}_{g,n})$ [Gor14] for $g>1$ in terms of the orbifold calculation of Harer and Zagier [HZ86]. Gorsky’s formula and Getzler’s results can in principle be passed through Getzler and Kapranov’s formula to determine $\chi^{S_{n}}(\overline{\mathcal{M}}_{g,n})$ for arbitrary $g$ and $n$, though this requires daunting calculations [Dia20].
@@
When it comes to Kontsevich moduli spaces, the most complete results on their topology and intersection theory are limited to genus zero. The $S_{n}$-equivariant Serre characteristics of both $\mathcal{M}_{0,n}(\mathbb{P}^{r},d)$ and ¯$\overline{\mathcal{M}}_{0,n}(\mathbb{P}^{r},d)$ were computed by Getzler and Pandharipande [GP06], by introducing symmetric functions in the stable map setting. Since ¯$\overline{\mathcal{M}}_{0,n}(\mathbb{P}^{r},d)$ is smooth and proper, its Serre characteristic determines its Hodge decomposition. Due to its role in enumerative geometry, there has been much interest in the intersection theory of ¯$\overline{\mathcal{M}}_{0,n}(\mathbb{P}^{r},d)$ [Pan99, BO03, BF03, Opr06a, Opr06b, MM07, MM08, KM10]. The rational cohomology of the interior is also well-understood [Pan98, FW16], as is the integral Chow ring for odd values of $d$ [CF22]. There has also been recent progress on the cohomology of moduli spaces of logarithmic stable maps to $\mathbb{P}^{r}$ in genus zero [Kan23, KHNSZ24, Dev24].
@@
In positive genus, most of the literature on the cohomology of $\mathcal{M}_{g,n}(\mathbb{P}^{r},d)$ and ¯$\overline{\mathcal{M}}_{g,n}(\mathbb{P}^{r},d)$ is focused on specific classes of interest in Gromov–Witten theory [GP99, FP00, FP03]. There are few known calculations that apply to the entire cohomology ring, likely due to the difficult singularities and excess components of the moduli space. Vakil and Zinger desingularised the so-called main component of ¯$\overline{\mathcal{M}}_{1,n}(\mathbb{P}^{r},d)$ [VZ08] to produce a smooth compactification ~$\widetilde{\mathcal{M}}_{1,n}(\mathbb{P}^{r},d)$ of $\mathcal{M}_{1,n}(\mathbb{P}^{r},d)$ with tractable geometry. Their work was later given a modular interpretation in logarithmic geometry by Ranganathan, Santos–Parker, and Wise [RSPW19] and extended to genus two by Battistella and Carocci [BC23]. We used the compactification ~$\widetilde{\mathcal{M}}_{1,n}(\mathbb{P}^{r},d)$ to compute the boundary complex of $\mathcal{M}_{1,n}(\mathbb{P}^{r},d)$ in [KS24], and we deduced the vanishing of the weight-zero compactly-supported cohomology $W_{0}H^{\star}_{c}(\mathcal{M}_{1,n}(\mathbb{P}^{r},d);\mathbb{Q})$. The second-named author computes the pure-weight rational cohomology $W_{\star}H^{\star}(\overline{\mathcal{M}}_{1,n}(\mathbb{P}^{r},d);\mathbb{Q})$ in forthcoming work [Son24]. Bae has recently initiated the study of tautological rings of Kontsevich spaces in positive genus [Bae20].
@@
Topologists have proven many interesting theorems about moduli spaces of maps, going back to Segal [Seg79]. Results in this direction usually focus on the space of maps from a fixed curve [Gue95, Ban22]. Tosteson [Tos22] has proven one of the few positive genus structural results about the rational cohomology of ¯$\overline{\mathcal{M}}_{g,n}(\mathbb{P}^{r},d)$ of which we are aware: for fixed $g,r,d$, and $i$, he proves that the sequence of $S_{n}$-representations determined by $H_{i}(\overline{\mathcal{M}}_{g,n}(\mathbb{P}^{r},d);\mathbb{Q})$ satisfies a form of representation stability, in the sense of Sam and Snowden [SS17]. His theorem implies several constraints on the asymptotic behaviour of these representations.
@@
On the combinatorial side, our technique of enumerating graphs with specified automorphisms by counting quotient graphs is inspired by the calculation of the $S_{n}$-equivariant top-weight Euler characteristic of $\mathcal{M}_{g,n}$ in [CFGP23]. Our method is reminiscent of Pólya enumeration [Pól37, PR87]. Adjacent problems on the enumeration of graph colourings and necklaces have been addressed [Mor72, Rob68, Han85]. Our use of wreath product symmetric functions is inspired by Petersen’s combinatorial proof [Pet12] of Getzler’s formula for $\mathsf{e}^{S_{n}}(\overline{\mathcal{M}}_{1,n})$ as well as his work on operad structures for moduli spaces of admissible $G$-covers [Pet13].
@@ 1.6. Further directions
It is possible that our approach to the topological Euler characteristics applies to the virtual Euler characteristics of ¯$\overline{\mathcal{M}}_{1,n}(\mathbb{P}^{r},d)$ [Giv00], [Lee04]. Among the key ingredients of our method, torus localisation [Giv15b, ACT20] and $S_{n}$-actions [Giv15a] have been explored in quantum K-theory, and the virtual Euler characteristic satisfies cut-and-paste properties relative to the forgetful morphism ¯$\overline{\mathcal{M}}_{1,n}(\mathbb{P}^{r},d)\to\mathfrak{M}_{1,n}$. Hence, there is reason to expect that analogous calculations could contribute to positive genus K-theoretic Gromov–Witten invariants.
@@
There are alternative modular compactifications of $\mathcal{M}_{1,n}(\mathbb{P}^{r},d)$, most notably the moduli space of quasimaps $\overline{\mathcal{Q}}_{1,n}(\mathbb{P}^{r},d)$ [MOP11] and the Vakil–Zinger desingularisation of ¯$\overline{\mathcal{M}}_{1,n}(\mathbb{P}^{r},d)$ [VZ08, RSPW19, KS24]. It would be interesting to compare the topology and combinatorics of these spaces. The locus ¯$\overline{\mathcal{M}}_{1}^{\mathrm{nrt}}(r)$ is common to all three compactifications, so one could hope for plethystic formulas comparing the Serre characteristics. We expect that our techniques could eventually give motivic analogues of wall-crossing formulas relating Gromov–Witten and quasimap invariants [CFK20], similar to the motivic wall-crossing formulas for Hassett spaces obtained in [KSY24].
Higher genus calculations
The Graber–Pandharipande description of ¯$\overline{\mathcal{M}}_{g,n}(\mathbb{P}^{r},d)^{\mathbb{C}^{\star}}$ implies that its Serre characteristic is determined by products of Serre characteristics of moduli spaces of stable curves, modulo finite groups. We expect that this should manifest in a formula, akin to the result [GK98, Theorem 8.13] discussed in §1.5, which expresses $\mathsf{e}(\overline{\mathcal{M}}_{g}(r)^{\mathbb{C}^{\star}})$ in terms of the series $\overline{\mathsf{m}}$ defined in (1.8), which we hope to pursue in future work. Our proof of Theorem A suggests that wreath product symmetric functions could be a useful technical tool for proving such a formula.
@@
The enumeration of coloured cycles in our work gives the count of $\mathbb{C}^{*}$-fixed strata in ¯$\overline{\mathcal{M}}_{1,n}^{\mathrm{nrt}}(\mathbb{P}^{r},d)$. It would be interesting if our techniques could be adapted to count all $\mathbb{C}^{\star}$-fixed strata in ¯$\overline{\mathcal{M}}_{1,n}(\mathbb{P}^{r},d)$. Such enumeration problems encode the combinatorial complexity of the compactification. In the case of stable curves $\mathcal{M}_{g,n}\subset\overline{\mathcal{M}}_{g,n}$, Aluffi–Marcolli–Nascimento [AMN24] enumerate the boundary strata of fixed dimension when $g=0$, and Getzler enumerates all boundary strata when $g=1$ [Get98]. To our knowledge, this direction has not been pursued for mapping spaces.
@@ Acknowledgements
We thank Dhruv Ranganathan and Burt Totaro for edifying conversations. We also thank Melody Chan, Megan Chang-Lee, Dan Petersen, Dhruv Ranganathan, and Burt Totaro for providing feedback on a draft of this article. SK is supported by NSF DMS-2401850, and he is grateful for the hospitality extended by members of the DPMMS at the University of Cambridge, where much of this work was carried out. TS is supported by a Cambridge Trust international scholarship.
@@ 2. Symmetric functions and wreath products
In this section we recall the background on symmetric functions and their wreath product analogues that we will use throughout the paper. We refer to [Mac95] and [GK98, §7] for more detailed discussions of this material. Given a partition $\lambda\vdash n$, we write $\lambda_{i}$ for the number of times $i$ appears in $\lambda$, so $\sum_{i=1}^{n}i\lambda_{i}=n$, and we write $|\lambda|:=n$. Given a permutation $\sigma\in S_{n}$, we write $\lambda(\sigma)\vdash n$ for the cycle type of $\sigma$.
@@ 2.1. Symmetric functions and $S_{n}$-equivariant characteristics
[…]
@@
for the Schur function corresponding to the partition $\lambda$. There is an inner product $\langle$-,$\rangle$ - , - ⟩, called the Hall inner product, encoding the inner product of characters of $S_{n}$. In particular, the Schur functions $s_{\lambda}$ for $\lambda\vdash n$ form an orthonormal basis for the homogeneous $\Lambda_{n}$ with respect to this inner product. For an arbitrary $S_{n}$-representation $V$, we have that
@@
i.e. the inner product of $\operatorname{ch}_{n}(V)$ with $s_{\lambda}$ is given by the multiplicity of the irreducible $V_{\lambda}$ in $V$. Now suppose that $V$ is an $S_{n}$-representation in the abelian category $\mathsf{M}$ of mixed Hodge structures over $\mathbb{Q}$. The Peter–Weyl theorem for representations of finite groups in $\mathsf{M}$ [Get95b, Theorem 3.2] implies that for each $\lambda\vdash n$, there exists a unique mixed Hodge structure $M_{\lambda}$ such that
@@
where $V_{\lambda}$ is a Specht module, considered as a mixed Hodge structure of type (0,0)(0,0)( 0 , 0 ). We then set
@@
and the element $\mathsf{e}_{\lambda}(X)$ should be thought of the virtual multiplicity of the Specht module $V_{\lambda}$ in $\mathsf{e}^{S_{n}}(X)$. There is a homomorphism
@@
[…]
2.2. $\mathbb{S}$-modules
@@
(2.9) $[\mathcal{V}]\mapsto\operatorname{ch}(\mathcal{V}):=\sum_{n\geq 0}\mathrm{ch}_{n}^{\mathsf{M}}(\mathcal{V}(n))\in K_{0}(\mathsf{M})\otimes\Lambda,$

### v2

@@ 1.3. Enumerative combinatorics of cycles
$\phantom{space}+\sum_{d\geq 2}q^{d}\sum_{k=2}^{d}\left(\eta_{k,d}(r)\frac{(1+2\dot{\mathscr{A}}_{0})^{2}}{(1-\psi_{2}(\mathscr{A}_{0}^{\prime\prime}))^{k/2+1}}+\sum_{j\mid k}\theta_{j,k,d}(r)\frac{1}{(1-\psi_{j}(\mathscr{A}_{0}^{\prime\prime}))^{k/j}}\right).$
The Serre characteristic of $\mathsf{Dih}_{r}$ is the solution to a graph enumeration problem: it is a count of isomorphism classes of localisation graphs, where each graph $\mathcal{G}$ is weighted by the type-$B$ character of the action of $\operatorname{Aut}(\mathcal{G})$ on $H(\mathcal{G})$. We can write this as
@@ 1.4. History and context
We are inspired by work of Getzler [22, 23] and Getzler–Kapranov [21] from the 1990s, using symmetric functions to work systematically with $S_{n}$-equivariant graph stratifications of moduli spaces of stable curves. Getzler and Kapranov work in the setting of modular operads, and give a formula [21, Theorem 8.13] which encodes how the two generating functions
@@
in $K_{0}(\mathsf{M})\otimes\Lambda[[q]]$ determine one another, in terms of plethysm and other algebraic operations on $\Lambda[[q]]$. Getzler computed $\mathsf{m}$ for $g\leq 1$ geometrically, and he then used the symmetric function formalism to compute $\overline{\mathsf{m}}$ for $g\leq 1$ [22, 23, 24]. In higher genus, the framework developed by Getzler and Kapranov has become a standard tool for passing cohomological information between moduli spaces of smooth and stable curves: see [9, §2] or [53, §2.5]. Gorsky has computed the $S_{n}$-equivariant Euler characteristic $\chi^{S_{n}}(\mathcal{M}_{g,n})$ [30] for $g>1$ in terms of the orbifold calculation of Harer and Zagier [34]. Gorsky’s formula and Getzler’s results can in principle be passed through Getzler and Kapranov’s formula to determine $\chi^{S_{n}}(\overline{\mathcal{M}}_{g,n})$ for arbitrary $g$ and $n$, though this requires daunting calculations [17].
@@
When it comes to Kontsevich moduli spaces, the most complete results on their topology and intersection theory are limited to genus zero. The $S_{n}$-equivariant Serre characteristics of both $\mathcal{M}_{0,n}(\mathbb{P}^{r},d)$ and $\overline{\mathcal{M}}_{0,n}(\mathbb{P}^{r},d)$ were computed by Getzler and Pandharipande [25], by introducing symmetric functions in the stable map setting. Since $\overline{\mathcal{M}}_{0,n}(\mathbb{P}^{r},d)$ is smooth and proper, its Serre characteristic determines its Hodge decomposition. Due to its role in enumerative geometry, there has been much interest in the intersection theory of $\overline{\mathcal{M}}_{0,n}(\mathbb{P}^{r},d)$ [52, 6, 10, 49, 50, 48, 47, 40]. The rational cohomology of the interior is also well-understood [51, 20], as is the integral Chow ring for odd values of $d$ [12]. There has also been recent progress on the cohomology of moduli spaces of logarithmic stable maps to $\mathbb{P}^{r}$ in genus zero [38, 39, 16].
@@
In positive genus, most of the literature on the cohomology of $\mathcal{M}_{g,n}(\mathbb{P}^{r},d)$ and $\overline{\mathcal{M}}_{g,n}(\mathbb{P}^{r},d)$ is focused on specific classes of interest in Gromov–Witten theory [31, 18, 19]. There are few known calculations that apply to the entire cohomology ring, likely due to the difficult singularities and excess components of the moduli space. Vakil and Zinger desingularised the so-called main component of $\overline{\mathcal{M}}_{1,n}(\mathbb{P}^{r},d)$ [66] to produce a smooth compactification $\widetilde{\mathcal{M}}_{1,n}(\mathbb{P}^{r},d)$ of $\mathcal{M}_{1,n}(\mathbb{P}^{r},d)$ with tractable geometry. Their work was later given a modular interpretation in logarithmic geometry by Ranganathan, Santos–Parker, and Wise [58] and extended to genus two by Battistella and Carocci [5]. We used the compactification $\widetilde{\mathcal{M}}_{1,n}(\mathbb{P}^{r},d)$ to compute the boundary complex of $\mathcal{M}_{1,n}(\mathbb{P}^{r},d)$ in [37], and we deduced the vanishing of the weight-zero compactly-supported cohomology $W_{0}H^{\star}_{c}(\mathcal{M}_{1,n}(\mathbb{P}^{r},d);\mathbb{Q})$. The second-named author computes the pure-weight rational cohomology $W_{\star}H^{\star}(\overline{\mathcal{M}}_{1,n}(\mathbb{P}^{r},d);\mathbb{Q})$ in [62]. Bae has recently initiated the study of tautological rings of Kontsevich spaces in positive genus [3].
@@
Topologists have proven many interesting theorems about moduli spaces of maps, going back to Segal [61]. Results in this direction usually focus on the space of maps from a fixed curve [32, 4]. Tosteson [65] has proven one of the few positive genus structural results about the rational cohomology of $\overline{\mathcal{M}}_{g,n}(\mathbb{P}^{r},d)$ of which we are aware: for fixed $g,r,d$, and $i$, he proves that the sequence of $S_{n}$-representations determined by $H_{i}(\overline{\mathcal{M}}_{g,n}(\mathbb{P}^{r},d);\mathbb{Q})$ satisfies a form of representation stability, in the sense of Sam and Snowden [60]. His theorem implies several constraints on the asymptotic behaviour of these representations.
@@
On the combinatorial side, our technique of enumerating graphs with specified automorphisms by counting quotient graphs is inspired by the calculation of the $S_{n}$-equivariant top-weight Euler characteristic of $\mathcal{M}_{g,n}$ in [13]. Our method is reminiscent of Pólya enumeration [57, 56]. Adjacent problems on the enumeration of graph colourings and necklaces have been addressed [46, 59, 33]. Our use of wreath product symmetric functions is inspired by work of Macdonald [44], Petersen’s combinatorial proof [54] of Getzler’s formula for $\mathsf{e}^{S_{n}}(\overline{\mathcal{M}}_{1,n})$, and his work on operad structures for moduli spaces of admissible $G$-covers [55].
@@ 1.5. Further directions
It is possible that our approach to the topological Euler characteristics applies to the virtual Euler characteristics of $\overline{\mathcal{M}}_{1,n}(\mathbb{P}^{r},d)$ [27], [42]. Among the key ingredients of our method, torus localisation [29, 2] and $S_{n}$-actions [28] have been explored in quantum K-theory, and the virtual Euler characteristic satisfies cut-and-paste properties relative to the forgetful morphism $\overline{\mathcal{M}}_{1,n}(\mathbb{P}^{r},d)\to\mathfrak{M}_{1,n}$. Hence, there is reason to expect that analogous calculations could contribute to positive genus K-theoretic Gromov–Witten invariants.
Higher genus calculations
The Graber–Pandharipande description of $\overline{\mathcal{M}}_{g,n}(\mathbb{P}^{r},d)^{\mathbb{C}^{\star}}$ implies that its Serre characteristic is determined by products of Serre characteristics of moduli spaces of stable curves, modulo finite groups. In [36], we prove a graph sum formula expressing the generating function $\sum_{g,n}\chi^{S_{n}}(\overline{\mathcal{M}}_{g,n}(\mathbb{P}^{r},d))t^{g-1}$ in terms of the generating function $\overline{\mathsf{m}}$ from (1.7) and certain graph sums. The present work calculates the $t$-degree 0 term by computing the genus one graph sums explicitly after separating rational tails. Algorithms to compute the graph sums in higher genus would make many more calculations possible.
@@
There are alternative modular compactifications of $\mathcal{M}_{1,n}(\mathbb{P}^{r},d)$, most notably the moduli space of quasimaps $\overline{\mathcal{Q}}_{1,n}(\mathbb{P}^{r},d)$ [45] and the Vakil–Zinger desingularisation of $\overline{\mathcal{M}}_{1,n}(\mathbb{P}^{r},d)$ [66, 58, 37]. It would be interesting to compare the topology and combinatorics of these spaces. The locus $\overline{\mathcal{M}}_{1}^{\mathrm{nrt}}(r)$ is common to all three compactifications, so one could hope for plethystic formulas comparing the Serre characteristics. We expect that our techniques could eventually give motivic analogues of wall-crossing formulas relating Gromov–Witten and quasimap invariants [14], similar to the motivic wall-crossing formulas for Hassett spaces obtained in [35].
@@
The enumeration of coloured cycles in our work gives the count of $\mathbb{C}^{*}$-fixed strata in $\overline{\mathcal{M}}_{1,n}^{\mathrm{nrt}}(\mathbb{P}^{r},d)$. It would be interesting if our techniques could be adapted to count all $\mathbb{C}^{\star}$-fixed strata in $\overline{\mathcal{M}}_{1,n}(\mathbb{P}^{r},d)$. Such enumeration problems encode the combinatorial complexity of the compactification. In the case of stable curves $\mathcal{M}_{g,n}\subset\overline{\mathcal{M}}_{g,n}$, Aluffi–Marcolli–Nascimento [1] enumerate the boundary strata of fixed dimension when $g=0$, and Getzler enumerates all boundary strata when $g=1$ [23]. To our knowledge, this direction has not been pursued for mapping spaces.
@@ Acknowledgements
We thank Dhruv Ranganathan and Burt Totaro for edifying conversations. We also thank Sarah Brauner, Melody Chan, Megan Chang-Lee, Dan Petersen, Dhruv Ranganathan, and Burt Totaro for providing feedback on a draft of this article. We would like to acknowledge helpful comments from anonymous referees. SK is supported by NSF DMS-2401850, and he is grateful for the hospitality extended by members of the DPMMS at the University of Cambridge, where much of this work was carried out. TS is supported by a Cambridge Trust international scholarship.
@@ 2. Symmetric functions and wreath products
In this section we recall the background on symmetric functions and their wreath product analogues that we will use throughout the paper. We refer to [44, 43] and [21, §7] for more detailed discussions of this material. Given a partition $\lambda\vdash n$, we write $\lambda_{i}$ for the number of times $i$ appears in $\lambda$, so $\sum_{i=1}^{n}i\lambda_{i}=n$, and we write $|\lambda|:=n$. Given a permutation $\sigma\in S_{n}$, we write $\lambda(\sigma)\vdash n$ for the cycle type of $\sigma$.
@@ 2.1. Symmetric functions and $S_{n}$-equivariant characteristics
It is convenient to identify $p_{0}=1$. We set $\Lambda_{n}\subset\Lambda$ for the homogeneous degree-$n$ part of $\Lambda$, so $p_{n}\in\Lambda_{n}$ for all $n\geq 0$. If $V$ is an $S_{n}$-representation in the category of $\mathbb{Q}$-vector spaces, we define its Frobenius characteristic by
@@
for the Schur function corresponding to the partition $\lambda$. There is an inner product $\langle-,$-$\rangle$, called the Hall inner product, encoding the inner product of characters of $S_{n}$. In particular, the Schur functions $s_{\lambda}$ for $\lambda\vdash n$ form an orthonormal basis for the homogeneous $\Lambda_{n}$ with respect to this inner product. For an arbitrary $S_{n}$-representation $V$, we have that
@@
i.e. the inner product of $\operatorname{ch}_{n}(V)$ with $s_{\lambda}$ is given by the multiplicity of the irreducible $V_{\lambda}$ in $V$. Now suppose that $V$ is an $S_{n}$-representation in the abelian category $\mathsf{M}$ of mixed Hodge structures over $\mathbb{Q}$. The Peter–Weyl theorem for representations of finite groups in $\mathsf{M}$ [26, Theorem 3.2] implies that for each $\lambda\vdash n$, there exists a unique mixed Hodge structure $M_{\lambda}$ such that
@@
where $V_{\lambda}$ is a Specht module, considered as a mixed Hodge structure of type $(0,0)$. We then set
@@
and the element $\mathsf{e}_{\lambda}(X)$ should be thought of as the virtual multiplicity of the Specht module $V_{\lambda}$ in $\mathsf{e}^{S_{n}}(X)$. There is a homomorphism
@@
Plethysm extends to $K_{0}(\mathsf{M})\otimes\Lambda$ by setting $h_{n}\circ[V]=[\mathrm{Sym}^{n}(V)]$ for $[V]$ the class of a mixed Hodge structure $V$.
$2.$ $\mathbb{S}$-modules
@@
(2.9) $\mathcal{V}]\mapsto\operatorname{ch}(\mathcal{V}):=\sum_{n\geq 0}\mathrm{ch}_{n}^{\mathsf{M}}(\mathcal{V}(n))\in K_{0}(\mathsf{M})\otimes\Lambda,$